Consider the function h = {(-3,4),(-2,2),(-1,0),(0,1),(1,3),(2,4),(3,-1} and function k={(-3,-2),(-2,0),(-1,-4),(0,0),(1,-3),(2,1),(3,2)}.
Compute the following expression if it exists: (h/k)(-2)
What in the world is going on here!

Answer by greenestamps(13200)



The notation means the only thing it could possibly mean: (h/k)(x) = h(x)/k(x). So

(h/k)(-2) = h(-2)/k(-2) = 2/0

Since division by 0 is not allowed, the answer is that (h/k)(-2) does NOT exist.

As further examples, note that (h/k)(0) would also not exist, because k(0)=0 and again division by 0 is not allowed.

But (h/k) exists for all the other possible input values; for example, (h/k)(-3) = h(-3)/k(-3) = 4/-2 = -2.